Problem 3 A charge Q is placed at the centre of the square of side 2.10 cm, at the corners of which four identical charges q-7.5 C are placed. Find the value of the charge Q so that the whole system is in equilibrium cmtacd n te value of the charge Oa! 7.98 What is the condition on the net force on a charge for it to be in equilibrium? Subonit AnswerIncorrect. Tries 1/6 Previous Trics
Answer #1

Individual forces.

Forces from the adjacent corners sum to
F1 = 2k(q/s)2/sqrt(2) = sqrt(2)k(q/s)2
Force from the opposite corner, where distance = sqrt(2)s,
F2 = k(q/(sqrt(2)s))2 = 0.5k(q/s)2
Total force F = F1+F2 = (sqrt(2)+0.5)k(q/s)2
Now just calculate -F = kQq/r2 ==> Q = -Fr2/(kq), where r = s/sqrt(2), or
Q = -Fs2/(2kq) = (sqrt(2)+0.5)q/2.
Note this is an unstable equilibrium.
For the values given, I get Q = -7.178 C